OpenStudy (anonymous):

Evaluate m^0 - n^2 for m = 2 and n = -1. a) -1 b) 0 c) 1

OpenStudy (anonymous):

b

OpenStudy (anonymous):

2^0=1 and (-1)^2=1 therefore 1-1=0